List of Publications

Francesco Baldassarri

[1] “Una funzione intera p-adica a valori interi”, Ann. Sc. Norm. Sup. Pisa, Classe di Scienze, Ser. IV, Vol.II, 1975, pp.321-331;

[2] “On second order linear differential equations with algebraic solutions”, with B. Dwork, Amer. J. of Math., Vol. 101, 1979, pp. 42-75;

[3] “On second order linear differential equations with algebraic solutions on algebraic curves”, Amer. J. of Math., Vol. 102, 1980, pp. 517-535;

[4] “On algebraic solutions of Lamé’s differential equation”, J. of D.E., Vol.41, No. 1, 1981, pp. 44-58;

[5] “Differential modules and singular points of p-adic differential equations”, Advances in Math., Vol.44, No. 2, 1982, pp. 155-179;

[6] “On inseparable descent”, Ann. Sc. Norm. Sup. Pisa, Classe di Scienze, Ser. IV, Vol. IX, 1982, pp.443-462;

[7] “Higher p-adic gamma functions and Dwork cohomology”, Soc. Math. de France, Astérisque, Vol.119-120, 1984, pp. 111-127;

[8] “Cohomologie p-adique pour la fonction _{3}F_{2}( \begin{matrix}a, & b_{1} & , b_{2} \\ & c_{1} & c_{2} \end{matrix} ; λ)”, Soc. Math.

de France, Astérisque, Vol.119-120, 1984, pp. 51-100;

[9] “Soluzioni algebriche dell’equazione di Lamé e torsione delle curve ellittiche”, Rend. Sem. Mat. e Fis. Milano, Vol.LVII (1987);

[10] “ Comparaison entre la cohomologie algébrique et la cohomologie p-adique rigide à coefficients dans un module différentiel I” , Inv. Math. 87 (1987), 83-99;

[11] “ Comparaison entre la cohomologie algébrique et la cohomologie p-adique rigide à coefficients dans un module différentiel II”, Math. Ann. 280 (1988), 417-439;

[12] “On singular projective structures”, with B.Dwork and F. Tovena, J. of D.E., Vol.80, 1989, pp.364-378;

[13] “p-adic Analysis”, Proceedings of the International Conference held in Trento, Italy May 29-June 3, 1989, Springer Lecture Notes in Mathematics, vol. 1454;

[14] “p-adic interpolation of Evans sums and deformation of the Selberg integrals: an application of Dwork’s theory”, Proceedings of the Katata Workshop on Special Differential Equations, Aug.26-31, 1991; [15] “Algebraic versus rigid cohomology with logarithmic coefficients”, Proceed-

ings of the Kyoto Symposium on Special Differential Equations, Sept.2-5, 1991, RIMS Publications;

[16] “ On Christol’s theorem. A generalization to systems of PDE’s with logarithmic singularities depending upon parameters”, with B.Chiarellotto, in p-adic Methods in Number Theory and Algebraic Geometry, A.Adolphson, S.Sperber, M.Tretkoff Eds., Contemporary Mathematics, Volume 133, AMS Publications 1992, 1-24;

[17] “ On André’s transfer theorem”, with B.Chiarellotto, in p-adic Methods in Number Theory and Algebraic Geometry, A.Adolphson, S.Sperber, M.Tretkoff Eds., Contemporary Mathematics, Volume 133, AMS Publications 1992, 25-37;

[18] “ Formal and p-adic theory of differential systems with logarithmic singularities depending upon parameters”, with B.Chiarellotto, Duke Math. J. , 72 (1993), p. 241-300;

[19] “ Algebraic versus rigid cohomology with logarithmic coefficients”, with B.Chiarellotto, in Barsotti Memorial Symposium, Perspectives in Mathematics Vol. 15, Academic Press 1994, p.11-50;

[20] “Special values of symmetric hypergeometric functions”, Trans. of the AMS, 348 (1996), p. 2249-2289;

[21] André Y., Baldassarri, F. : “Geometric theory of G-functions”, in “Arithmetic Geometry”, F. Catanese Ed., Cambridge University Press 1997.

[22] ”On arithmetic size of linear differential equations”, with L. Di Vizio, in preparation;

[23] “p-adic formulas and unit root sub-crystals of the hypergeometric system”, with M. Cailotto, in preparation.

[24] “Notes on differential modules I. Regularity in several variables”, with Y. André, in preparation;

[25] “Notes on differential modules II. Irregularity in several variables”, with Y. André, in preparation;

[26] “Notes on differential modules III. Direct images (the Gauss-Manin connection)”, with Y. André, in preparation;

[27] “Notes on differential modules IV. Complex and p-adic comparison theorems”, with Y. André, in preparation.

Dec. 12, 1997